A calorimeter contains 18.0 mL of water at 13.0 ∘C . When 1.80 g of X (a substance with a molar mass of 82.0 g/mol ) is added, it dissolves via the reaction

X(s)+H2O(l)→X(aq)

and the temperature of the solution increases to 27.5 ∘C .

Calculate the enthalpy change, ΔH, for this reaction per mole of X.

Assume that the specific heat of the resulting solution is equal to that of water [4.18 J/(g⋅∘C)], that density of water is 1.00 g/mL, and that no heat is lost to the calorimeter itself, nor to the surroundings.

Express the change in enthalpy in kilojoules per mole to three significant figures.
Answer #1

Density of water = 1.00 g/ml

Total mass = Mass of water + Mass of substance = 18.0 ml * 1.00 g/ml + 1.80 = 19.80 [Assuming that it dissolves via the reaction and added to weight, change the mass if this assumption is incorrect]

Heat provided = Total mass * specific heat capacity * Change in temperature = 19.80 g * 4.18 J/gC * (27.5-13.0)C = 1200.078 J

Enthalpy Change = -Heat provided/(number of moles) = -1200.078/(1.80/82.0) = -54.7 kJ/mol

Note - post any doubts/queries in comments section.